OpenStudy (anonymous):

Factor: 36 -(y - 4)^2

OpenStudy (anonymous):

first take out the GCF

OpenStudy (anonymous):

Difference of squares.\[\ \ \ \ 36-(y-4)^2\\=(6+(y-4))(6-(y-4))\\=-(y+2)(y-10)\]

OpenStudy (anonymous):

how did you get the last part?

OpenStudy (anonymous):

@v_ejc \[(6+(y-4))(6-(y-4))=(6+y-4)(6-y+4)=(y+2)(10-y)=-(y+2)(y-10)\]